OpenStudy (anonymous):

4.6x10 to the power of 2=

OpenStudy (anonymous):

hello numbers!

OpenStudy (anonymous):

Hii

OpenStudy (anonymous):

\[4.6\times 10^2=4.6\times 100=460\]

OpenStudy (saifoo.khan):

Long time no C!

OpenStudy (anonymous):

long time no C++